Reports of 40419
Detector Operation (Operations Report)
ruggi - 23:41 Tuesday 19 February 2019 (44937) Print this report
Comment to Operator report - Afternoon shift (44935)

I leave the IMC AA closed in full bandwidth, after the implementation of a sensing matrix which looks fine, and more important after a change of the working point, which rely on the zero of the FF and NF quadrants. The eventual improvement of the system is entirely due to the work done by optics group for the quality of the signals.

AdV-TCS (CO2 laser projector)
nardecchia, rocchi, carbognani - 23:02 Tuesday 19 February 2019 (44936) Print this report
Comment to Rotation Stage replacement on the CO2 benches (44909)
Today we installed all the 8 half-wave plates in the new rotation stages on the CO2 benches.
Tomorrow we will complete the check of the proper operation of the rotation stages and we will set the right powers for all the actuation lines.
We turned on both lasers at 19.26 UTC to do the power tuning with the lasers at the steady state (fig. 1).
We left both CO2 lasers ON with the beams blocked on the benches.

NB: In the VPM the new process TCSAgilisRot had software problem and did not restarted, to be checked by Franco.
Images attached to this comment
Detector Operation (Operations Report)
berni - 22:59 Tuesday 19 February 2019 (44935) Print this report
Operator report - Afternoon shift
The main activity of this shift was the work on SDB2 in DET LAB; people kept working up to 20:30 UTC.
In parallel other activities were carried out:
- ISYS activities; still in progress
- TCS CO2 lasers switching ON; completed at 20:00 UTC.
- work at NE building by Polish group; completed at 18:30 UTC.

At around 20:00 UTC I was contacted by Romain because he wanted to perform some activities with the ITF locked in PR_ITF; unfortunately I had no instructions to lock the ITF in that state without B1p and I could reach only LOCKED_RECOMBINED. We decided to postpone the activity.

I left Paolo still working on the ISYS.

SAT
We mirror coils left unplugged

SBE
At 16:48 UTC I closed the SNEB LC; they had been left open by Alessandro.

ISYS
yesterday night (around 4:00 UTC) we had a jump of 1 degree in the temperatures of PSL_AMP_DIODE_4; see attached pictures.
Images attached to this report
Comments to this report:
ruggi - 23:41 Tuesday 19 February 2019 (44937) Print this report

I leave the IMC AA closed in full bandwidth, after the implementation of a sensing matrix which looks fine, and more important after a change of the working point, which rely on the zero of the FF and NF quadrants. The eventual improvement of the system is entirely due to the work done by optics group for the quality of the signals.

Environmental Monitoring (Environmental Monitoring)
fiori, paoletti - 21:22 Tuesday 19 February 2019 (44934) Print this report
ENV_EIB_ACC_Z changed to ENV_EIB_ACC_Y

The name has been changed (see title) to correctly account for the orientation of the accelerometer, which is vertically oriented.

Environmental Monitoring (Environmental Monitoring)
fiori - 21:16 Tuesday 19 February 2019 (44933) Print this report
magnetic probe

Today I positioned the magnetic probe ENV_TCS_MAG next to CO2 lasers interlock devices in TCS electronics room (see picture).

Images attached to this report
Environmental Monitoring (Environmental Monitoring)
fiori, menendez, romero - 18:01 Tuesday 19 February 2019 (44931) Print this report
Comment to Seismic and acoustic injections (44824)
Today we have computed the noise projection of the acoustic injections in the laser labs for different auxiliary channels, among which we have ENV_EIB_MIC and BsX_TX. We have found out that the noise projection corresponding to ENV_EIB_MIC, BsX_TX and BsX_TY make sense, given that the projection is below the quiet noise. However, in the case of BsX_X and BsX_Y we have seen that at 622 Hz, the projection goes above the quiet time, which is something we don't quite understand. Still, in order to check whether the program was computing everything correctly, we computed the noise projection for 622 Hz by hand . For that purpose, we looked for the following data:
1) Value of Hrec during injection at the given frequency: 7.65e-23 (1/sqrt(Hz))
2) Value of BsX_Y during injection at the given frequency: 9.5e-04 (1/sqrt(Hz))
3) Value of BsX_Y during quiet time at the given frequency: 6.5e-04 (1/sqrt(Hz))
With this we could compute the transfer function at that frequency, which resulted in: 8e-20, which is precisely what the program calculated. Then, the noise projection ends up being: 5.2e-23 (1/sqrt(Hz)), which was also in agreement with the program's output.
Images attached to this comment
AdV-COM (AdV commissioning (1st part) )
casanueva - 17:35 Tuesday 19 February 2019 (44930) Print this report
Unlocks during the mini-run

I have analyzed the unlocks of the mini-run, from Saturday at 20.45 UTC to this monday morning. I have found 23 unlocks.

  • 9 unlocks while on LN3 :
    • 2 due to glitches on the LCs of the IMC
    • 7 due to Fast Unlocks
  • 14 unlocks during lock acquisition:
    • 3 during the engagement of the boost -> Looked like due to low gain (Figure 1)
    • 1 due to some glitches ~125Hz that saturated the controls of MICH and DARM at MICH 0.2 -> Could be a problem of high gain on DARM, since it happened while the gain was still increasing (Figure 2)
    • 1 during the alignment of the PR -> the target position was wrong
    • 9 due to Fast unlocks
Images attached to this report
AdV-INJ (General opto-mechanical layout, common parts construction, Installation and pre-commissioning)
Paoletti, Montanari - 16:19 Tuesday 19 February 2019 (44928) Print this report
Change of DCD120D HV power supply with upgraded version
Yesterday Monday Feb. 18th we have changed the "in use" DCD120D HV regulators with an upgraded version having a lower noise and a fold-back current limit circuit (100mA max).
The two units are locate in the IER (aka Piscina) and in the Atrium (the room between the SAS and the Laser Lab).
Computers Network (Storage)
salconi - 16:16 Tuesday 19 February 2019 (44929) Print this report
Circular buffer deletion: v31 v32 on Feb 21th
on Feb, 21th the following rawdata volumes will be deleted in order to make space for new files:

v31
v32

that will involve data from:
GPS: 1220869100 13 Sep 2018 10:18:02 UTC (13 Sep 2018 12:18:02 CEST)
to:
GPS: 1221315900 18 Sep 2018 14:24:42 UTC (18 Sep 2018 16:24:42 CEST)
Detector Operation (Operations Report)
Montanari - 16:02 Tuesday 19 February 2019 (44927) Print this report
Operator Report - Morning Shift Feb 19

This morning I found the ITF locked in LOCKED ARMS.

The morning was dedicated to the maintanance activities, started at 08:10UTC::
- HVAC: check of the boiler at terminal building
- INF: standard Cleaning of the Experimental Buildings
- VAC: continued the activity about SR-BS turbo pump replacement
           standard vacuum maintenance and refill
- tiltometer movement to NE building
- TCS: chiller check and refill, both in particular NI
- SBE: SDB2 work
- INJ: cable installation in the Atrium by electronics
         working on the injection alignment

DMS
As requested by the TCS expert, I changed the thresholds about the CO2 Laser temperature monitoring

AdV-DAQ (Calibration)
bonnand, rolland - 15:36 Tuesday 19 February 2019 (44924) Print this report
Measurement for B1 and B1p timing calibration

Profiting of SDB2 opening this week, Romain has enlighted the B1 and B1p photodiodes with the LED flashing the IRIG-B coming from a GPS receiver from the DAQ-room via a 20 m cable. As a reminder, the channels B1_PD2_DC,Audio and B1p_PD1_DC,Audio are now acquired at 1 MHz permanently with the DS suffix. The channels have been copied from the raw_full data and stored in /data/procdata/calibration/rawdataCalibration/2019_02_photodiodeTiming/

  • B1_PD2: 1234618800, for 10 s
  • B1p_PD1: 1234618990, for 10 s
  • B1p_PD2:  1234619160, for 10 s
  • B1_PD1: 1234619220, for 10 s

 

AdV-PSL (Pre-mode-cleaner block)
cleva - 15:25 Tuesday 19 February 2019 (44926) Print this report
Comment to Offset at the slave and PMC error signal & neoVAN alignment (43914)

- V_PZT_Monit is indeed 1/20 V_PZT

it means the calibration of the slow piezo mentionned on plot 1 is 46 MHz/V instead

- please replace plot 2 with the attced plot (gives the foreseen zoom)

 

 

Images attached to this comment
AdV-PSL (Laser Bench setup)
cleva - 15:13 Tuesday 19 February 2019 (44925) Print this report
Comment to Slave & PMC check, PZT TF (44601)

From plot there is only 3,5 dB between the pzt channels (instead of the 10 dB mentionned)

Images attached to this comment
Detector Characterisation (Glitches)
mwas - 12:58 Tuesday 19 February 2019 (44923) Print this report
Comment to Unknown glitch family at 31 Hz (44910)
Figure 1. I have looked at the spectra during one of the loudest glitches, and it has a 1/f^3 shape. This is the frequency shape of a step function (1/f), and then filtered by the mirror suspension (1/f^2).
Figure 2, 3 and 4 are examples of some the loudest glitches. There is clearly a step in the correction sent to the mirrors, at least for figure 2.
In the three examples the glitch correspond to a power drop on B1_PD2 and then an oscillation caused by DARM loop . It would be intersting if the glitch always start with B1_PD2 power going down.
Note that the answer might be different for the very loud glitches (SNR>100), and for the more moderate glitches.
Images attached to this comment
Injection system (EIB-SAS)
fiori, derossi, paoletti - 12:41 Tuesday 19 February 2019 (44922) Print this report
Comment to Accelerometer 50Hz line (44470)

After having investigated the accelerometer conditioning electronics in EERoom and found nothing, today we investigated the part on EIB.

We found the culprint was the low-noise coaxial 10-32 to BNC cable that happened to be damaged maybe during the last intervention.

Camilla replaced the cable, and the accelerometer now looks ok (Figure 1, purple is now).

Images attached to this comment
Detector Characterisation (Tools)
hemming - 11:13 Tuesday 19 February 2019 (44920) Print this report
VIM update pushed to production

Last Friday, subsequent to the adaptation of VIM for federated authentication, the latest version of the software was pushed to production. Information relating to the modified version is available here.

The web interface of VIM will iterate over a few different versions over the next period, as various features requested by users are integrated into the software. These upgrades will be reported in the Logbook, but please let me know if you notice anything functioning incorrectly.

Note also that this report has been added in the newly-created Detector Characterisation -> Tools task.

Detector Operation (Operations Report)
mwas - 9:34 Tuesday 19 February 2019 (44921) Print this report
Comment to Operator Report - Afternoon Shift 17 Febrary 2019 (44901)
I have run the noise budget at the time of the range drop.
Figure 1 shows it is an issue with one of the calibration injections
Figure 2 shows it is the WE photon calibrator.
The calibration team will look into why it happened, the pcal was running for the whole weekend, but something strange must have happened during these 30 minutes.

Images attached to this comment
Detector Operation (Operations Report)
Gherardini - 22:55 Monday 18 February 2019 (44528) Print this report
Operator report - afternoon shift 18/2
Several activity had been carried out in this afternoon:

- SDB2 work, in progress...
- injection system alignment (#44917).
- WE suspension electronic check.
- rotator stage installation on CO2 bench (#44909), in progress...
- SR-BS turbo pump replacement, in progress...

to not loose the mirrors alignment I relocked the two cavities and left the itf in LOCKED_ARMS state.

-DAQ
at 18:31utc the TprogImgDET, DET_Img_CEB server crashed and the rtpc11 became unreachable, expert informed and problem fixed.

-INF
- UTA INJ in "Portata Nominale" from 14:15UTC to 19:25UTC
- UTA DET in "Portata Nominale" from 9:00 UTC and it is still running in this configuration
AdV-PSL (Laser Bench setup)
Cleva, Coulon - 22:34 Monday 18 February 2019 (44919) Print this report
Comment to Slave & PMC check, PZT TF (44601)

Images attached to this comment
Injection system (General activities)
Cleva, Coulon - 22:05 Monday 18 February 2019 (44918) Print this report
Comment to videodriver swapped again (44912)

Unless wrong, removing the 50 ohms attached to the amplifier which follows the Videodriver increases the gain after the videodriver by a factor 2 while keeping the dynamic of that amplifier compliant.

Hence 0.5V at videodriver output becomes 1V once the 50 ohm is removed, and then 100 V at the following amplifier output, with no saturation at those levels. This would release the pressure on the videodriver by a factor 2. The bandwidth should not be affected.

 

Simple to check since the 50 ohm is a BNC one attached to the input of the amplifier on the LB close to the crystal

 

 

Injection system (General activities)
derossi, pillant, gosselin, sposito, coulon, chiummo, paoletti - 20:33 Monday 18 February 2019 (44912) Print this report
videodriver swapped again

This morning we swapped again the videodriver and we installed a modified one with a new filter, whose zero is at 3 kHz (it was at 14 Hz before), in order to reduce the rms by avoiding to send EOM corrections at 50, 100 and 150 Hz. Its trasfer function is shown in the first plot.

The IMC is currently locking with an attenuation from 5 to 14 dB (until 7 dB it shows a loop oscillation @ 270 kHz), and we left it with 15 dB. Since by increasing the attenuation from 7 to 15 dB we lowered the general gain of the rampeauto, we then increased the EOM corrections thanks to the videodriver gain. The overall effect should be to release the piezo corrections.

With this videodriver saturation is now occurring at 0.5 V rather than 0.4 V (an example is given in the 2nd plot), so that we gained a little bit in the actuation range.

We finally remeasured the IMC trasfer function, which is shown on the 3rd plot in blue, compared with the previous one in pink.

 

Images attached to this report
Comments to this report:
Cleva, Coulon - 22:05 Monday 18 February 2019 (44918) Print this report

Unless wrong, removing the 50 ohms attached to the amplifier which follows the Videodriver increases the gain after the videodriver by a factor 2 while keeping the dynamic of that amplifier compliant.

Hence 0.5V at videodriver output becomes 1V once the 50 ohm is removed, and then 100 V at the following amplifier output, with no saturation at those levels. This would release the pressure on the videodriver by a factor 2. The bandwidth should not be affected.

 

Simple to check since the 50 ohm is a BNC one attached to the input of the amplifier on the LB close to the crystal

 

 

Injection system (Input beam alignment - hardware and software)
gosselin, derossi, pillant - 20:12 Monday 18 February 2019 (44917) Print this report
Beam realignement on EIB
This afternoon we did a general realignment of the beam going through the EIB, from the laser bench to the IMC.
We started from the first modulator (6MHz 8 MHz) where it appeared to be miscentered (see figure 1 and 2, before and after realignment). Then we went through the other modulators, the Faraday Isolator, the IPC and recentered the beam on the alignment/matching mirrors.
Then we finely aligned the beam to bring the correction of the BPC close to 0 (figure 6). The loop closed normally.

This misalignment on the first modulator must explain the beam shape of the IMC reflection that we had before (figure 3) with respect to the one that we have now (figure 4).
Now it looks like we have pure mismatch. We measured it to be 5.7% (figure 5) instead of the about 9% that we had before.

We then slightly realigned the IMC cavity and measured the new sensing matrix of the Automatic Alignment. The signals are now much cleaner. We closed the loop in drift control.
Images attached to this report
AdV-TCS (CO2 laser projector)
nardecchia, nenci, sposito, carbognani - 18:56 Monday 18 February 2019 (44909) Print this report
Rotation Stage replacement on the CO2 benches
MICOS rotation stages used on the CO2 benches to change the actuation power are not properly working, indeed some of them rotate only in one direction and when they reached the end-stops the controller needs to be manually reset.
Thus, in view of O3, we decided to replace the MICOS ones with the Newport Agilis rotation stages (already used in INJ).
We turned off both CO2 lasers at 09.19 UTC in order to allow to the electronic people to change the connectors of the cables between the TCS room and the benches.
Then Franco worked on the software in order to implement the remote control.
Tomorrow we will continue with the installation.

Comments to this report:
nardecchia, rocchi, carbognani - 23:02 Tuesday 19 February 2019 (44936) Print this report
Today we installed all the 8 half-wave plates in the new rotation stages on the CO2 benches.
Tomorrow we will complete the check of the proper operation of the rotation stages and we will set the right powers for all the actuation lines.
We turned on both lasers at 19.26 UTC to do the power tuning with the lasers at the steady state (fig. 1).
We left both CO2 lasers ON with the beams blocked on the benches.

NB: In the VPM the new process TCSAgilisRot had software problem and did not restarted, to be checked by Franco.
Images attached to this comment
Detector Operation (Operations Report)
swinkels - 18:50 Monday 18 February 2019 (44914) Print this report
Comment to Operator Report - Afternoon Shift 17 Febrary 2019 (44901)
I did a brute-force correlation of the half-hour long drop in the range with all channels in trend.ffl. Other than some expected correlation with some DARM BRMS and Hrec related channels, I could not find any smoking gun.
AdV-COM (AdV commissioning (1st part) )
ruggi - 17:12 Monday 18 February 2019 (44915) Print this report
Peaks around 50 Hz

The spectrum of DARM around 50 Hz seems to be composed by several small structures. It is known that at ~ 48 Hz there is a vertical resonance of FP payloads (involving the F7 crossbar). In order to investigate if some of those resonances are present in DARM, a noise injection has been performed.

Fig 1: noise has been injected on NI_MAR_Y_CORR and WI_MAR_Y CORR, at two different times. In figure, also the correction currently applied by the vertical damper is shown.

Fig 2: the effect on DARM of WI noise injection. The expected resonance is visible, but the projection of damper correction, obtained by this measurement, is quite low.

Fig 3: the same consideration is valid for NI noise injection. In this case the frequency of the crossbar resonance (48.3 Hz, already measured by the local sensor) is close, but not equal, to the frequency of one peak in the clean data (48.4 Hz). I would exclude that the NI crossbar resonance is the source of that peak. But the noise injection has excited one more peak (47.8 Hz), the origin of which is not known. It would be better to check if this structure can be more easily excited acting on the marionette in a different direction.

The work has to be continued injecting noise on NE and WE.

 

Images attached to this report
Search
×

Warning

×